OpenStudy (anonymous):

Rationalize the numerator. (√28)/7

OpenStudy (jdoe0001):

\(\bf \cfrac{\sqrt{a}}{b} \times \cfrac{\sqrt{a}}{\sqrt{a}}\)

OpenStudy (anonymous):

is the answer